OpenStudy (anonymous):

Suppose f(x)=4x-2 and g(x)=2x+1..... Find the value of f(-3)/g(-2)

OpenStudy (anonymous):

in f(x) plug -3 in for x, and in g(x) plug in -2 for x so youd have f(-3)=(4(-3))-2=-14 g(-2)=(2(-2))+1=-3 f(-3)/g(-2)=-14/-3=14/3

OpenStudy (anonymous):

Ohhh thank you.

OpenStudy (anonymous):

anytime =)